### v1

# CrashFixer: A crash resolution agent for the Linux kernel

Abstract.
Code large language models (LLMs) have shown impressive capabilities on a multitude of software engineering tasks. In particular, they have demonstrated remarkable utility in the task of code repair. However, common benchmarks used to evaluate the performance of code LLMs are often limited to small-scale settings. In this work, we build upon kGym (Mathai et al., [n. d.]), which shares a benchmark for system-level Linux kernel bugs and a platform to run experiments on the Linux kernel.
This paper introduces CrashFixer, the first LLM-based software repair agent that is applicable to Linux kernel bugs. Inspired by the typical workflow of a kernel developer, we identify the key capabilities an expert developer leverages to resolve a kernel crash. Using this as our guide, we revisit the kGym platform and identify key system improvements needed to practically run LLM-based agents at the scale of the Linux kernel ($\sim 50K$ files and $\sim 20M$ lines of code). We implement these changes by extending kGym to create an improved platform – called kGymSuite, which will be open-sourced. Finally, the paper presents an evaluation of various repair strategies for such complex kernel bugs and showcases the value of explicitly generating a hypothesis before attempting to fix bugs in complex systems such as the Linux kernel. We also evaluated CrashFixer’s capabilities on still open bugs, and found at least two patch suggestions considered plausible to resolve the reported bug.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; ; isbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Large Language Models (LLMs) have shown impressive capabilities across a wide range of software engineering tasks, including code generation (Chen et al., 2021), test creation (Kang et al., 2023), automated code review (Li et al., 2022), and program repair(Gao et al., 2022)—debugging and generating candidate patches for known failures. While most of the prior work targets application-level software (Yang et al., 2024; Zhang et al., 2024a), we aim to take a significant step toward the more complex challenge of automatically repairing Linux-kernel (lin, 2025) crashes as reported by fuzzers (Miller et al., 1990). The Syzkaller fuzzer (syz, 2025) has exposed over $5,000$ bugs in the kernel to date and continues to find more, yet resolving these crashes remains a difficult and time-intensive task. Given the kernel codebase, a crash report, and the crash triggering input, our goal is to patch the code so that the input no longer causes a crash, while also maintaining system functionality.
Repairing kernel crashes is inherently challenging for LLMs, as the available inputs are typically crash dumps (e.g., stack traces, kernel memory dumps, register contents, etc.) rather than natural language (NL) bug reports or curated unit test cases. NL descriptions are generally more accessible for both humans and LLMs to interpret, while unit test cases provide concrete guidance for reproducing and understanding failures. In contrast, our focus—automatically resolving kernel crashes discovered via fuzzing—is considerably more demanding. The Linux kernel is a vast, low-level, concurrent codebase written in C and assembly, with privileged hardware access and stringent coding rules. Fuzzers identify fail-stop crashes without accompanying human-readable explanations or correctness specifications, limiting the effectiveness of standard LLM-based repair approaches. Nevertheless, fuzzers do collect structured crash reports and sanitizer logs, offering an opportunity to automate the entire crash resolution pipeline: reproduction, diagnosis, and patching.
To address this challenge, we introduce CrashFixer, the first LLM-based software agent designed to repair Linux kernel crashes that are discovered through fuzzing. Unlike prior approaches that attempt a single-shot resolution of kernel failures (Mathai et al., [n. d.]), CrashFixer closely emulates a kernel developer’s iterative and tool-assisted debugging process to resolve kernel crashes. With the crash context and source code in hand, our key insight is to follow the actions of a kernel developer: first, interpret the available execution traces and relevant code to form an understanding of the failure, and then, generate the patch based on that understanding. Following this, CrashFixer works in two phases:
  1. (1)
    Hypothesis Generation. It begins by invoking low-level diagnostic tools to reproduce the crash and collect minimal yet relevant execution traces. These traces help isolate the context and conditions under which the crash may occur. With the crash context and source code in hand, CrashFixer constructs a natural language hypothesis describing the likely root cause of the crash and a plausible mitigation plan by leveraging the LLM’s reasoning capabilities.
  2. (2)
    Patch Generation. Based on this hypothesis, it synthesizes a diverse set of candidate patches. CrashFixer then filters out patches that fail to compile and chooses the one best aligned with the hypothesized mitigation plan. CrashFixer proceeds to apply this patch. After applying the patch, it re-runs the crash-inducing input to validate whether the failure has been resolved. If the crash persists or if new issues are introduced by the patch, CrashFixer uses the observed behavior to refine its next hypothesis.
To improve chances of success, CrashFixer explores parallel trajectories across both phases, with each trajectory leading to a different diagnostic path and set of corrective actions, forming a forest-like structure of patch exploration. This allows CrashFixer to successfully arrive at a patch in the case of a good exploration cycle, while mitigating the impacts of imprecise patches/hypothesis in unsuccessful explorations. Thus, CrashFixer can robustly navigate the vast space of potential fixes while maintaining the flexibility of pursuing promising hypotheses or abandoning poor hypotheses during this search.
To support the CrashFixer workflow, we develop kGymSuite, a significantly enhanced version of kGym (Mathai et al., [n. d.]), serving both as a runtime system and a comprehensive experimental validation platform. kGymSuite integrates a suite of efficient and essential tools tailored to the challenges of kernel crash diagnosis and repair. It introduces intelligent mechanisms to identify and extract minimal yet informative execution traces, even in the presence of disruptive kernel crashes. These traces are critical for providing actionable context to CrashFixer ’s reasoning process. Additionally, kGymSuite includes substantial improvements in build caching, crash reproduction fidelity, and patch validation speed—optimizations that are vital for the iterative nature of LLM-based debugging. These enhancements not only power CrashFixer ’s capabilities but also establish kGymSuite as a reusable infrastructure for the broader research community. By enabling reliable, scalable, and reproducible evaluations of automated kernel repair techniques, kGymSuite lays the groundwork for advancing LLM-driven systems-level debugging.
We have deployed CrashFixer and kGymSuite to Google Cloud and conducted over 500,000 kernel patching experiments. We evaluate CrashFixer’s capabilities on both kBenchSyz —a benchmark of curated Syzkaller bugs (Mathai et al., [n. d.])—and recent open Syzkaller bugs. On kBenchSyz, CrashFixer resolves $\sim 49\%$ of kernel crashes automatically with file-level localization assistance (i.e., when the file containing the developer fix is disclosed to CrashFixer). Providing a minimal, relevant execution trace preceding the crash, results in a relative performance increase of $\sim 10\%$. In an ongoing study of applying CrashFixer to resolve the latest unaddressed Syzkaller bugs , we find two plausible patches and continue to investigate other plausible solutions. A detailed manual analysis of the CrashFixer patches for $68$ kBenchSyz bugs shows that $(1)$ for $10$ bugs, CrashFixer generates accurate natural language diagnoses and nearly identical patches to the ground-truth developers’ fixes; $(2)$ for $9$ bugs, the generated diagnoses and patches effectively help pinpoint root causes; and $(3)$ for the remaining $49$ bugs, CrashFixer successfully (but perhaps questionably) resolves the crashes, but through “amputation” (i.e. bypassing faulty code all together) or through significant modification of existing functionality. This raises an interesting research question: how to guide LLMs to preserve intended defined behaviors of code when resolving undefined behaviors such as crashes, which we leave for future work.
This paper makes the following key contributions:
(1) CrashFixer, the first LLM-based agent for resolving Linux kernel crashes found by fuzzing. It introduces a novel inference pipeline that mirrors how developers investigate kernel crashes step by step, and leverages minimal, relevant execution history to improve crash resolution. To the best of our knowledge, CrashFixer is the first execution-aware program repair agent for the Linux Kernel. The design principles behind CrashFixer have the potential to generalize to agents resolving fuzzing-induced crashes in other kernels and user-space applications.
(2) kGymSuite, a substantially enhanced kernel debugging and experimental validation platform based on kGym. It captures targeted execution traces even in the presence of kernel crashes and improves performance and scalability. kGymSuite can support future LLM-based program repair research in the kernel domain.
(3) An evaluation of one of the leading LLMs (Gemini Pro 1.5) in CrashFixer, demonstrating its ability to autonomously resolve—or assist a human in resolving—a substantial fraction of Linux kernel crashes in a public benchmark. Some patches modify (or even amputate) both buggy and functional code, which is an open problem that will require future research to address.
The impact of our work is considerable: kernel developers are tragically overworked and the number of open bugs increases faster than the number of fixes. Providing actionable complete (or even partial) patches with human-understandable rationales at even a modest true-positive rate can help rebalance this losing battle by increasing patch-fix velocity and improving kernel security and reliability.
2. Related Work
LLM Abilities and Software Engineering (SE). Recent advancements have enabled LLMs to show code comprehension and completion capabilities (lla, 2025; Rozière et al., 2024; Nijkamp et al., 2023; Fried et al., 2023; Chen et al., 2021), leading to their inclusion in various SE tasks like testing (Xia et al., 2024b; Wang et al., 2024; Kang et al., 2023), program repair (Dinh et al., 2023; Gao et al., 2022; Rondon et al., 2025), commit generation (Liu et al., 2023), and pull request reviews (Li et al., 2022). In this paper, we focus on program repair. We note, that unlike our work, previous literature has not attempted program repair for large systems-level repositories like Linux. Through CrashFixer, we hope to explore the current capabilities of LLMs in this domain.
Benchmarks. The most commonly evaluated application of Code LLMs is code generation. As a result, there are a plethora of code completion benchmarks. Most benchmarks such as HumanEval (Chen et al., 2021), primarily assess code completion at a file level, though recent efforts have introduced cross-file (Shrivastava et al., 2023; Pei et al., 2023; Zhang et al., 2023; Ding et al., 2023), repository-level (Jimenez et al., 2024), and kernel-level benchmarks like kBenchSyz and kBenchC (Mathai et al., [n. d.]). As CrashFixer targets Linux kernel crashes, we choose to measure our efficacy on kBenchSyz, a curated benchmark of Syzkaller bugs. Each kBenchSyz bug has associated metadata like the git commit id, the kernel configuration file, the reproducer that triggers the bug, the developer fix, the crash report, and auxiliary information such as developer discussions.
Execution History. Previous works have explored the use of execution information at runtime to enhance LLM-generated code. Some approaches rely on costly manual annotation and prompt engineering (Chen et al., 2023b; Wu et al., 2023), which, while effective, are not scalable for large-scale SE. Other methods refine code generation by incorporating additional contextual information, such as interpreter outputs and self-generated explanations, alongside execution data (Chen et al., 2018, 2023a; Hu et al., 2024). Some pipelines train models to refine a base model’s output (Pearce et al., 2023; Gupta et al., 2023), whereas others rely on the base model itself (Le et al., 2022; Shinn et al., 2023; Zhong et al., 2024). CrashFixer follows the latter approach, where we try to improve performance by analyzing execution logs relevant to a kernel crash. However, it differs fundamentally in scope, as previous work has largely focused on file-level benchmarks, whereas CrashFixer extends execution history analysis to a broader, system-level setting. This distinction makes CrashFixer the first to systematically investigate the impact of execution traces on debugging at the scale of the Linux kernel.
LLMs for SE Agents. LLMs have also been used for end-to-end software engineering tasks in agentic systems, where one or more models autonomously plan and execute actions by invoking tools and observing feedback from the environment. In the domain of SE, agents are often used for automated program repair where, given some buggy code, the agent makes decisions to ultimately generate an appropriate patch. Previous agents have been equipped with various tools for code search, code generation, patch application, and test execution (Yang et al., 2024; Arora et al., 2024; Zhang et al., 2024a; Ruan et al., 2024; Liu et al., 2024; Rondon et al., 2025). Most of these agents, however, have been designed to resolve user/developer reported issues similar to SWE-bench (Jimenez et al., 2024); these code issues are from popular open-source Python libraries with reliable natural language descriptions of the task. Recently, Passerine (Rondon et al., 2025) has addressed the more challenging class of multi-lingual user-space bugs at Google, including sanitizer-found bugs. In contrast, CrashFixer is designed to resolve crashes in the Linux kernel, a large repository with concurrent, low-level code and sanitizer-generated crash reports.
kGym. A key platform we reference in this work is kGym (Mathai et al., [n. d.]), designed for research on Linux kernel crash resolution. In particular, kGym is a platform that can consistently reproduce kernel crashes and evaluate if a given fix resolves a kernel crash. To do so kGym provides two key functionalities. The first is a kernel-building pipeline to compile various kernels when given relevant specifications. The second is a crash-reproduction pipeline that takes a pre-built kernel and a crash-triggering program (or “reproducer”) and attempts to reliably reproduce a kernel crash by parallelly executing the reproducer on multiple virtual machines. Despite its many applications, we identify bottlenecks in kGym that noticeably affect kernel build times and key missing features like the inability to collect detailed execution traces when running the reproducer. Such issues are remedied in our enhanced implementation kGymSuite. For user-space automated program repair, a few different “gym”-like environments are available, including SWE-Bench itself and, recently, debug-gym (Yuan et al., 2025).
Figure 1. The crash-fixing workflow of a kernel developer. CrashFixer follows this flow to generate hypothesis and patch.
Localization Candidates During crash resolution, we narrow down a small set of source files likely responsible for the crash. They are called localization candidates.
Hypothesis A hypothesis is a NL explanation that identifies a plausible root cause of the crash and proposes a mitigation strategy.
Patch A patch is a candidate fix that modifies the Linux kernel source code. In this paper, we assume that patches only modify the localization candidates.
Reproducer Run Each Syzkaller bug includes a reproducer program that, when executed on the compiled kernel, triggers the crash. We define a reproducer run as executing this program for up to 10 minutes or until a crash occurs.
Crash Resolution The goal of the crash-fixing process is to generate a patch that prevents the reproducer from triggering a crash. We define crash resolution as the task of generating syntactically valid and semantically plausible patches that eliminate the crash during a reproducer run.
Self Reflection Self-reflection is the process of asking an LLM to review its previously generated outputs. This is used to either modify its previous generation or, in the case of ranking, to select the best generated output.
Self Consistency Self-consistency is the process of asking an LLM to (1) first describe a plan, (2) suggest implementations of this plan, and (3) choose the implementation that is most consistent with the described plan.
Execution Logs During a reproducer run, the kernel executes a sequence of functions—some of which are defined in the localization candidates. We refer to this sequence as the (function-granularity) execution logs
Table 1. Terminologies used in CrashFixer and kGymSuite frameworks.
3. Motivation and Overview
This section begins by describing how kernel developers typically fix crashes (§3.1), which motivates the design of our system. It then presents an overview of the CrashFixer pipeline (§3.2) and highlights key improvements in kGymSuite that enable and support CrashFixer.
3.1. Developer Crash-Resolution Workflow
Fuzzers like Syzkaller offer an advantage of automating the discovery of certain bug classes. They may detect errors that are either fail-stop (e.g., null-pointer dereferences and segment violation errors) or have well-known failure detectors (e.g., sanitizer detectors for buffer overflows and use-after-free bugs, deadlock detectors, etc.). This allows developers to trade computational resources for the human effort of meticulous code inspection. Syzbot is a large-scale continuous fuzzing infrastructure that runs Syzkaller on the tips of the Linux kernel development trees.
When a new crash is found, Figure 1 illustrates a model of how a kernel developer typically resolves the crash, using a real bug reported by Syzkaller (syz, 2025). First, an automatic notification (See Figure 1: 1) is sent to the kernel developer mailing list. This report includes the bug type as identified by the sanitizer (e.g., “memory leak”), the function where the crash occurs (“dvb_create_media_graph”), the commit id of the kernel version, the build configuration, and scripts to reproduce the bug.
The developer then reviews the crash report to understand the nature of the issue. Because the crash was discovered via fuzzing, the report lacks a natural language description of the root cause. Instead, it contains machine-generated diagnostics such as the sanitizer-reported bug type and the kernel stack trace at the time of the crash ( 2). Based on their understanding of the codebase and the report, the developer identifies a few relevant source files to examine, formulates a hypothesis about the root cause ( 3), and modifies the code accordingly ( 4).
Next, the developer rebuilds the kernel and runs ( 5) the provided reproducer script to check whether the crash has been resolved. In many cases, the initial patch does not fully fix the issue ( 6). The developer then inspects kernel execution logs ( 7) to refine their understanding of the faulty code paths and iterates on the patch. This cycle of inspection, hypothesis refinement, and patching continues until the crash is resolved and the fix passes code review by kernel maintainers.
As seen from this workflow, kernel developers rely on three key capabilities for effective crash resolution, which directly inform our system design:
$(1)$ Iterative hypothesis testing. Developers iteratively formulate and refine hypotheses by editing the codebase and validating patches. Hypothesis testing helps developers converge on high-quality fixes through trial and error. Our system needs to follow the same principle, enabling LLM agents to iteratively improve patch quality. Moreover, LLM agents, lacking the deep contextual understanding of experienced developers, may require significantly more hypothesis-testing iterations. Our system must support this rapid experimentation loop at scale.
$(2)$ Execution inspection. Access to kernel execution logs is crucial for diagnosing crashes. However, a single reproducer run can generate a large volume of logs, easily exceeding input limits and reasoning capacities of current LLMs. Upon inspecting the logs for sample kBenchSyz bugs, we observe more than $10$M called functions during a single reproducer run. Thus, our system must extract minimal yet relevant execution traces to help agents reason effectively.
$(3)$ Efficient builds. Fast kernel builds are essential to test patches quickly. This becomes even more important for LLM agents, which often need to run many reproducer sessions across multiple patch attempts to find a valid fix.
Following these insights, Figure 2 illustrates our kernel crash resolution pipeline, which consists of two main components: (a) CrashFixer, our crash-resolution agent (Section 3.2); and (b) kGymSuite, our enhanced experimentation platform (Section 3.3). Table 1 presents the terminologies we used to describe our system.
3.2. CrashFixer Pipeline Overview
Figure 2. CrashFixer Pipeline.
The CrashFixer pipeline departs significantly from prior program repair efforts. The state-of-the-art system, kGym (Mathai et al., [n. d.]), attempts to fix kernel crashes in a single step—going directly from the crash report to patch generation. This one-shot approach exceeds the capabilities of current LLMs and, unsurprisingly, yields poor performance. Other generic coding agents (Dinh et al., 2023; Gao et al., 2022; Rondon et al., 2025) target user-space applications with NL descriptions of features or bugs. Their pipelines are typically relatively short: retrieve relevant code based on the issue description, then generate a patch using an LLM. In contrast, CrashFixer tackles a far more challenging problem and is, to the best of our knowledge, the first agent designed to resolve fuzzing-detected crashes and to incorporate execution logs into the repair process.
The pipeline begins by providing CrashFixer with the execution logs, crash report, and source code of the localization candidates (see Table 1 for their definitions). Using this input, CrashFixer generates a set of candidate hypotheses explaining the crash. It then performs self-reflection to select a plausible hypothesis and produces a set of patches implementing it. Before triggering expensive kernel builds, the patches undergo a fast compilation check to filter out syntactically invalid candidates. If multiple compilable patches remain, CrashFixer applies self-consistency reasoning to assess alignment with the selected hypothesis and chooses a single patch. Finally, CrashFixer uses kGymSuite to efficiently build the kernel, run the reproducer, and validate whether the crash is resolved. We detail each step of this pipeline in §4.
In our motivating example of memory leak bug (Figure 1), we select ‘dvbdev.c” as our localization candidate as it is in the execution log. CrashFixer then starts the crash resolution process and eventually generates a relevant hypothesis (green box in Figure 1). As shown, CrashFixer isolates the root cause of the crash (a missing kfree) and then suggests a mitigation plan. Using this, CrashFixer generates a robust patch depicted in Figure 1. The suggested patch is semantically equivalent to the corresponding developer fix.
3.3. kGymSuite Overview
To efficiently support the CrashFixer pipeline, the underlying platform must provide key system capabilities. However, the original kGym lacks the tooling needed to collect execution logs during reproducer runs. Since kernel execution generates a massive volume of logs (e.g., thousands of functions), easily overwhelming current LLMs, kGymSuite adds support to selectively collect and compress relevant execution traces (e.g., a dozen of functions) for use by CrashFixer5).
In addition, we identify a major bottleneck in kGym when validating patches and rebuilding the kernel repeatedly. kGym lacks a fast compilation check mechanism and is slow at building kernels, making large-scale experiments impractical. To address this, kGymSuite introduces cached kernel builds and efficient compilation checks (§5.2), dramatically reducing build times and enabling scalable crash resolution.
Next, we will present CrashFixer and kGymSuite in detail.
4. CrashFixer
Based on our preliminary experiments with LLMs to resolve crashes, we identify five challenges CrashFixer must address.
$(1)$ Decomposing Debug Cycle. Within a single debug cycle (i.e., debug-edit-run), going directly from crash report to correct code edit in one stage yields poor results. Even with retries, the patches often oscillate between vastly different edits. This single-stage approach also lacks interpretability: it provides no rationale for why a particular edit is proposed. As a result, the process is hard to steer or debug, especially when the LLM hallucinates and produces incorrect or irrelevant code changes. To address this, we decompose a debug cycle into two primary stages: Hypothesis (Section 4.1) & Patch generation (Section 4.2).
$(2)$ Hallucination. LLMs occasionally hallucinate, generating hypotheses or patches ungrounded in the input or inconsistent with the program context. For example, the model may propose a root cause that contradicts the execution trace or suggest edits to functions never invoked in the reproducer run. These hallucinations waste valuable debug attempts and can mislead the overall repair process. To mitigate this, CrashFixer must be internally consistent and anchor its reasoning to observable evidence, such as crash reports and execution logs. To this end, we deploy several strategies like self-consistency checking, self reflection, and ranking during hypothesis and patch generation.
$(3)$ Iterative Debug Cycle. LLM agents struggle to complete crash resolution in a single turn, regardless of how many internal stages that turn includes. This is expected—even kernel developers often require multiple attempts. For example, the crash report in Figure 1 points to the function dvb_create_media_graph (an allocator), which appears at the top of the call stack, naturally drawing initial developer attention. However, the true fix lies in the deallocator (dvb_media_device_free). Since LLMs lack deep kernel expertise, they too require multiple iterations of debug cycle to converge on a correct fix. This raises the question of how to organize these iterations so that later cycles build productively on earlier ones, while early mistakes do not derail subsequent progress, mirroring the classic exploration vs. exploitation trade-off. Section 4.3 illustrates our strategy.
$(4)$ Generated Patch should Compile. LLM-generated patches are often uncompilable. This is partly due to the specialized and underrepresented nature of kernel code in LLM training data. Another key reason is the use of the git diff format in kGym, which is sensitive to surrounding lines and line numbers. As acknowledged by prior work, this format causes many LLM-generated patches to be inapplicable (i.e., git apply fails), wasting valuable debug attempts.
$(5)$ Tool support for iterative patch generation. Resolving kernel crashes demands a different set of tradeoffs and tools compared to user-space program repair. For instance, building the Linux kernel and completing a reproducer run can take over 10 minutes, in stark contrast to the few seconds needed for an LLM call. As a result, CrashFixer must sometimes lean more heavily on the LLM to filter out unpromising patches before investing in expensive build and validation steps. Similarly, short execution traces preceding a crash often help kernel developers debug; thus, CrashFixer requires tooling to extract and provide such targeted traces.
In this section, we explain how CrashFixer tackles the first three challenges. Challenges $(4)$ and $(5)$ are addressed through the design and deployment of kGymSuite, detailed in Section 5. All prompts are provided in the appendix (see Appendix A.5 and Appendix A.6).
CrashFixer takes a crash report, execution log, and kernel source code as input, and tries to produce a patch that can repair the crash. Figure 2 shows an end-to-end pipeline of CrashFixer. It works in following three phases:
4.1. Hypothesis Generation & Selection
Given a kernel crash report, source code, and execution logs, we prompt CrashFixer to generate a set of plausible hypotheses that explain the root cause of the crash and propose corresponding mitigation strategies. However, LLMs are prone to hallucination—producing plausible-sounding but incorrect explanations based on fabricated or irrelevant information. This behavior has also been widely observed in tasks such as information retrieval (Huang et al., 2025).
To mitigate this, we incorporate a self-reflection step. After generating hypotheses, we ask CrashFixer to reflect on the full set by reviewing each one, identifying its strengths and weaknesses, and selecting the hypothesis most consistent with the observed crash. While this self-reflection process is not immune to hallucination, combining generation with structured critique and ranking can reduce error, similar to asking a human to approach the same problem from multiple angles to avoid systematic bias.
4.2. Patch Generation
After selecting a hypothesis, CrashFixer begins the patch generation process. To address the format issue with git diff patches, CrashFixer adopts two ideas.
Replace-based Edit Interfaces. Instead of requiring CrashFixer to produce a full git diff, we ask it to generate one or more edits as tuples of the form (original, replaced), where original is a snippet from the source code and replaced is the suggested replacement. Similar interfaces have been effective in other agent-based software repair settings (Zhang et al., 2024b; Xia et al., 2024a). Additionally, CrashFixer provides a rationale for each edit (edit reasoning), enabling users to trace and validate the intent behind the change. This is especially useful for reviewing patches that resolve crashes, allowing unnecessary or misguided edits to be filtered out.
Patch Selection. After selecting a hypothesis, CrashFixer generates multiple patch candidates using the replace-based interface. To filter the uncompilable ones, we invoke the compilation check tool in kGymSuite. As shown in Figure 2, kGymSuite uses the bug ID to load the corresponding SuiteCache (Section 5.2) and tests each patch for successful compilation.
After discarding uncompilable patches, we prompt CrashFixer to select the patch that best aligns with the selected hypothesis—a process called self-consistency . This step mimics a lightweight code review, where patches are evaluated based on their coherence with the intended fix. The selected patch is then passed to kGymSuite for a full kernel build.
If no candidate patches survive this stage, the selected hypothesis is deemed invalid, and CrashFixer restarts the process from hypothesis generation.
4.3. Iterative Debug & Repair
(a) $\mathbb{T}_{debug}$ ($\mathcal{D}=3$, $\mathcal{B}=2$)
(b) Example of $\mathbb{T}_{debug}$ exploration. In edit $1$, CrashFixer removes the original code inadvertently removing the first kfree but adds the missing kfree. In edit $2$, CrashFixer rectifies this by adding the original kfree.
Figure 3. Iterative Debug Cycle represented with a tree or forest. Each node $N_{i}$ represents a debug cycle and is associated with a state $(C_{i-1},R_{i-1})$—the codebase and resulting crash report. The tree is parameterized by maximum depth $\mathcal{D}$ and branching factor $\mathcal{B}$. The root node $N_{1}$ starts from $(C_{0},R_{0})$, the initial codebase and crash report.
To balance exploitation vs exploration, as mentioned in challenges $1$ $3$, CrashFixer organizes debug-cycles into either a tree (Figure 3(a)) or a forest (multiple parallel tree instances).
Debug Cycle. We define a debug-cycle as a single cycle encompassing the following steps - (a) hypothesis generation, (b) patch generation, (c) compiling the kernel, and (d) running the reproducer. In a debug-cycle, hypothesis generation can be optional depending on the step from which the CrashFixer tries alternative exploration.
Debug Tree. Each node in the tree corresponds to one debug cycle, and branching allows CrashFixer to explore alternative hypotheses when earlier attempts fail. We refer to this search structure as $\mathbb{T}_{debug}$. See Figure 3(a) for details. Each node $N_{i}$ represents a debug cycle and is associated with a state $(C_{i-1},R_{i-1})$ (the codebase and resulting crash report) from where the debug-cycle starts exploring. At any node, if an edit succeeds, exploration along that branch stops. If an edit fails to resolve the crash, child nodes (e.g., $N_{i+1}$, $N_{i+2}$) inherit the updated state $(C_{i},R_{i})$ and explore alternate hypotheses. This enables parallel search paths and incremental refinement across cycles. More advanced strategies (e.g., Monte Carlo Search (Zhou et al., 2024)) can be used to guide exploration.
Debug Forest. To mitigate error propagation from early hallucinated edits, we also consider a forest configuration: multiple shallow trees instead of one deep one. Thus, a forest $\mathcal{F}$ of two trees with depth $\mathcal{D}=3$ and $\mathcal{B}=1$ has roughly the same debug cycle budget as a single tree with $\mathcal{B}=2$, $\mathcal{D}=3$. Forests reduce the risk of compounding errors in long paths by enabling exploration from independent roots.
In Figure 3(b), we show an example of a tree search for the memory leak bug, where multiple debug-cycles help CrashFixer to get the correct patch.
5. kGymSuite
To support the CrashFixer pipeline, kGymSuite provides essential system and tooling support. First, it compresses relevant execution traces, avoiding the overload of full kernel logs (see Section 5.1). It also addresses performance bottlenecks in patch validation by introducing fast compilation checks and cached kernel builds, enabling scalable and efficient crash resolution (see Section 5.2).
5.1. Minimal Yet Relevant Execution Traces
Stack traces alone often miss the root cause—e.g., in Figure 1, the actual leak in dvb_media_device_free is not shown, while the trace points to the allocation site. Since kernel logs like dmesg are too noisy, kGymSuite uses precise execution tracing to help identify the true source of the bug.
Execution tracing is essential for debugging kernel crashes, but integrating it into our framework presents key challenges: (1) traces can span millions of tokens—far exceeding LLM context limits; (2) collecting complete traces is slow and especially difficult when a crash occurs; and (3) only a small fraction of the trace is typically relevant, turning the task into a needle-in-a-haystack problem. To address these issues, we design our tracing system around three goals: (i) efficient trace collection, (ii) resilience to crashes, and (iii) focusing only on trace segments most relevant to crash resolution. We achieve efficiency through in-memory function-level tracing, limited to files identified as localization candidates (Section 5.1.1). For crash tolerance, we extract traces from kernel crash-dump buffers after a failure (Section 5.1.2). To reduce irrelevant data, we isolate the crashing thread and extract trace segments around the execution of candidate functions (Section 5.1.3). We discuss these design decisions and discarded alternatives in the sections that follow.
5.1.1. Trace Collection
LLMs excel at understanding local code context, such as a few functions. Therefore, our trace-collection mechanism is designed to guide CrashFixer to the right functions, not to capture detailed control or data flow. We collect function-level execution traces from the kernel, allowing CrashFixer to focus on the relevant source code.
To achieve this, we implement a macro-based logging tool that writes to an in-memory ring buffer using Linux trace_printk tracing utility (Rostedt, [n. d.]). We wrap this macro to log the process ID and a per-file counter for ordering trace entries. We insert this macro at the start of every function for each file identified as a localization candidate. For instance, in our running example, since the localization candidate is dvbdev.c, all of its functions including dvb_create _media_graph and dvb_media_device_free are instrumented, even though not all are relevant to the specific bug.
Alternatives Considered and Rejected. We considered using printk for logging, but it causes high I/O overhead due to frequent file-system flushes, leading to hangs in  26.16% of kBenchSyz examples. We also evaluated ftrace, which avoids I/O issues and supports per-function tracing without code edits, but misses inlined or optimized-away functions (during compilation) —limiting CrashFixer ’s visibility and affecting some kBenchSyz data points.
5.1.2. Trace Retrieval
During interactive debugging, trace data from ftrace and trace_printk can be accessed via kernel APIs like /sys/kernel/debug/tracing/trace. However, these interfaces are unavailable after a crash, and serializing in-memory traces to logs is unreliable due to their large size. Instead, we use the kDump facility (kdu, 2025) to capture a full kernel memory snapshot at crash time, including the trace buffer. This snapshot is saved by a dump-capture kernel and extracted using crash-trace (fuj, 2025), followed by parsing with trace-cmd. We modify Syzkaller so that after a crash, kGymSuite connects via SSH to the dump-capture kernel and retrieves the kDump. This pipeline successfully recovers traces for 255 out of 279 kBenchSyz bugs (within 5 retries). The remaining 24 fail due to kDump misconfigurations or unsupported Debian VM image.
Alternatives Considered and Rejected. While Syzkaller keeps track of printk logs through VM serial port output, its performance overhead makes it impractical. Similarly, ftrace offers a dump_on_oops option that dumps in-memory traces via printk when the kernel crashes, but large trace sizes—up to 1.4 GB in our tests—lead to long logging times and dropped records due to excessive I/O.
5.1.3. Trace Minimization
Our trace minimization follows two key principles: (1) isolate the crashing thread from all active kernel threads, and (2) extract a focused portion of the function trace around the crash-report stack trace and the localized candidate file. This excludes bugs involving multi-thread interactions or cross-file logic, which we leave for future work. Given our focus on reproducible, single-thread bugs—most of which involve a single file—this is a reasonable scope.
The minimization algorithm (Figure 4) filters trace records for the crashing thread, anchors to the stack trace, and scans backward to match stack frames in order. It then expands forward and backward up to a limit (200 functions) or until a pattern repeats. For example, in our memory-leak case, it starts from PID 3062, anchors to dvb_create_media_graph, and expands to include relevant, traced functions, including dvb_media_device_free, which is the actual fix location. An example of the exact execution trace for the memory leak bug is shown in Appendix A.3.
Figure 4. The trace-minimization algorithm starts by scanning the execution trace backward to find all stack-trace functions (green) from the localized file in order, ignoring untraced ones (blue). This matching region, labeled “match,” may also include additional relevant functions like q and repeated calls like m. We refer to this first step as “stack trace anchoring”. Then, the span is expanded backwards and forwards until a repeating pattern of trace functions is detected (q, q, ... in the prefix, and p, m, p, m in the suffix), up to a maximum number of function records.
Although simple and effective, our minimization algorithm has some limitations. First, repeating patterns may demonstrate useful behavior in specific crash-reproduction scenarios (e.g., a crash that takes exactly 5 copies of a system call to manifest). Some reproducers from Syzkaller explicitly mention the need for such repetitions to trigger a crash, and this information can be used to avoid missing important parts of the execution. Second, although we match one possible point in the execution trace where an offending invocation occurred, there may be others; in our running example, multiple allocations of memory may have occurred due to dvd_create_media_graph, and we only identify the last one. For particularly complex bugs, it may be necessary for a future incarnation of our agent to pose additional hypotheses as to which of the matches of the stack trace may be the one to investigate.
5.1.4. Empirical Post-Rationalization
In what follows, we provide empirical evidence to support each step of our minimization algorithm by analyzing the collected executions for all bugs in kBenchSyz. We evaluate our minimization algorithm using the Complete Intersection Score (CIS), which measures how often all patch-edited functions (i.e. functions modified in the developer-fix) appear in a minimized trace—scoring 1 if complete, 0 otherwise, summed across all kBenchSyz bugs. Starting with only traced stack-trace functions yields a CIS of 104. Stack anchoring improves this to 108, and backward/forward expansion raises it to 117 and 127, respectively—our highest score. Including the full trace of the PID gives a CIS of 137 but exceeds LLM limits (2400 records vs. our average of 12). This analysis assumes ideal localization (i.e. localized file is the same as the developer-modified file) and focuses on single-file fixes, highlighting the value of minimization as one component of a larger pipeline.
5.2. Efficient Kernel Builds
Fast Kernel Builds. The Linux kernel is a massive system, comprising over $50K$ files and $20M$ lines of code—posing significant challenges for LLM-driven tooling. Building the kernel from scratch is slow, yet the prior platform kGym triggers a full rebuild for every patch attempt. On an 8-vCPU Google Compute Engine instance (c3-standard-8), this adds approximately 19 minutes per patching experiment.
In practice, however, kernel patches are typically small: developer-submitted fixes in our dataset modify just 1.28 files on average (and at most 7). Full rebuilds are thus wasteful.
kGymSuite addresses this with SuiteCache —a distributed kernel build cache shared across all LLM experiments. SuiteCache stores both the full source tree and precompiled object files for each kernel version in kBenchSyz. When a patch is applied, only the modified files are recompiled, and the resulting object files are linked with cached, unmodified ones. This optimization brings build time down to under 3 minutes on the same 8-vCPU GCE instance, including time spent downloading object files from the cache. Even when using compute-heavy 88-vCPU instances, we get an average speedup of 42.69% across the dataset (see Appendix A.1 for more details).
Fast Compilation Check. Since current LLMs often generate uncompilable patches, kGym frequently wastes time and resources setting up full kernel build jobs—cloning repositories, checking out source code, and applying backport commits—only to discover that the patched code fails to compile.
To avoid this, kGymSuite introduces a fast compilation check step prior to initiating a full build. A key observation is that most LLM-generated crash-resolution patches modify only .c files, not header files. In fact, all CrashFixer patches in our evaluation modify .c files. Therefore, it is sufficient to verify whether the modified .c files compile in isolation. kGymSuite does this using a targeted make command directed at each patched file. This check takes only 2–3 seconds per patch, allowing the system to quickly filter out uncompilable candidates and conserve resources.
6. Implementation and Evaluation
We have implemented and deployed CrashFixer and kGymSuite on Google Cloud Platform. The system runs on two clusters comprising 37 GCE machines each, with 88 vCPUs per machine, and has conducted over 500,000 LLM-driven patching and kernel build experiments to date.
Our evaluation is guided by the following questions:
  • RQ1: How effective is CrashFixer at resolving Linux kernel crashes? What is the contribution of each individual component in the pipeline?
  • RQ2: How useful are the CrashFixer-generated patches? Recall that successful crash resolution does not translate to correct patch.
  • RQ3: What is the cost and latency of CrashFixer?
6.1. Evaluation Setup
Datasets. We primarily evaluate CrashFixer using bugs from kBenchSyz, along with ongoing experiments on recent open Syzkaller bugs (Section 7). Following Mathai et al. ([n. d.]), we split kBenchSyz into two subsets: (a) the small context subset (SCS), containing 117 bugs, and (b) the large context subset (LCS), containing the remaining 162 bugs. Here, context refers to the number of tokens passed to the LLM. The localized files for each SCS bug fit within 16K tokens, while LCS bugs may require up to 50K tokens (Mathai et al., [n. d.]). This partitioning allows us to efficiently test strategies on SCS and apply the most effective ones to LCS.
Models. Our methodology is LLM-agnostic. For cost efficiency, we use Gemini 1.5 Pro (versions 001 and 002) as the underlying LLM. The results reported may not reflect the performance of newer Gemini models or LLMs from other providers.
Localization Candidates. For kBenchSyz bugs, we use the files modified by the developer fix to derive localization candidates, allowing us to better assess CrashFixer’s capabilities. For open bugs, we iteratively try each file referenced in the crash report, based on the observation that 74% of fixes in kBenchSyz modify only a single file in the crash report.
Strategy
Rea-
soning
[…]
Small Context Subset (117 bugs)
1. Baseline Top-10 (Mathai et al., [n. d.]) 10 1 1 10
2. Edit-Compile-Test 1 4 2 34.00 (32/36)
3. Edit-Compile-Test ++ 1 4 2 42.00 (42/42)
4. Hyp then Patch 1 4 2 46.33 (39/51)
5. Hyp-SR then Patch 1 4 2 48.33 (47/49)
6. Hyp-SR then Patch $\mathcal{F}$ 4 4 1 54.67 (51/57)
7. Hyp-SR then Patch-SC $\mathcal{F}$ 4 4 1 58.33 (57/59)
8. Hyp-SR then Patch-SC $\mathcal{F}$** 4 4 1 62.33 (60/64)
Long Context Subset (162 bugs)
9. Hyp-SR then Patch-SC $\mathcal{F}$ 4 4 1 65.33 (61/71)
10. Hyp-SR then Patch-SC $\mathcal{F}$** 4 4 1 75.00 (72/79)
Table 2. Average (min/max) number of crashes resolved in $3$ runs (SR: Self-reflection; SC: Self-consistency; $\mathcal{F}$: debug forests instead of debug tree; ** indicates Gemini-1.5-pro-002).
6.2. RQ1. Effectiveness
Table 2 (row $7$) shows that CrashFixer resolves $\sim 58$ bugs of SCS crashes—5.8$\times$ more than the simple agent in kGym. It also resolves $\sim 65$ bugs of LCS crashes (row $9$). With a stronger model (Gemini 1.5-pro-002 rows $8$ and $10$), mitigated crashes increase to $\sim 62$ (SCS) and $\sim 75$ (LCS). Table 2 ablates different CrashFixer settings as described below.
  • Edit-Compile-Test & Edit-Compile-Test++ (Table 2: rows 2 & 3). The basic Edit-Compile-Test loop resolves 34 Linux crashes by using (original, replaced) pairs without reasoning. Edit-Compile-Test++ improves this by adding edit-level explanations, resulting in a 23.52% relative performance gain.
  • Hypothesis Generation then Patch Generation (Table 2: rows 4 & 5). In this strategy, CrashFixer: (i) generates a hypothesis, (ii) creates candidate patches, (iii) filters out uncompilable ones, and (iv) builds and tests the first valid patch, eliminating failed builds. This resolves $\sim 46$ bugs, increasing to $\sim 48$ with self-reflection. CrashFixer uses temperature 0.8 for diverse hypotheses/patches, and 0.2 for other steps.
  • Introducing Forests (Table 2: rows 6–10). We see clear gains by decomposing the original tree into a forest of smaller, parallel $\mathbb{T}_{debug}$. Applying the Hyp then Patch strategy with 4 trees ($\mathcal{D}=4,\mathcal{B}=1$) raises performance to $\sim 54$ resolved bugs. This improvement stems from giving CrashFixer multiple fresh starting points, avoiding early edits that negatively impact downstream nodes. Adding self-consistency to select the patch best aligned with the generated hypothesis further boosts performance to $\sim 58$. This trend continues for the better model Gemini-1.5-pro-002 $\sim 62$ as well as long context bugs as well, up to $\sim 75$.
Adding Execution Traces. Only $255$ out of $279$ kBenchSyz bugs come with execution traces due to the reasons mentioned in Section 5.1.2. Thus, we evaluate the effectiveness of adding execution traces only on the $255$ bugs. In LCS, CrashFixer has a relative improvement of $10.36\%$ using 1.5-pro-002 and $13.61\%$ using 1.5-pro-001. In SCS, we see a relative increase of $9.7\%$ with 1.5-pro-002 and consistent results for 1.5-pro-001 (see Appendix A.2 for the detailed numbers).
Multi-turn. To understand the importance of multiple edit attempts, we measure the % of bugs only solved at depths greater than $1$ in SCS. For 1.5-pro-001, this accounts for $\sim 50\%$ of solved bugs in our final strategy (without execution) and $\sim 37.56\%$ with execution, demonstrating that tree-based exploration is broadly effective within the benchmark.
6.3. RQ2. Utility
Many plausible solutions can resolve a kernel crash, so predicting a patch equivalent to a developer’s is challenging. Sadly, a generated edit may resolve the kernel crash while compromising functionality. Thus, we manually analyze the patches for $68$ bugs in one run of Gemini 1.5-pro-002 (with execution) to understand the utility of the CrashFixer patches compared to the ground-truth fix patches for them.
Table 3 shows the utility results. We find that the patches fall into the following three categories. $(1)$ We consider patches as plausible if they are semantically equivalent to the developer patch. $10$ out of $68$ CrashFixer patches are plausible. $(2)$ We consider a patch as helpful if it targets the root cause correctly, but is not perfectly equivalent. $9$ CrashFixer patches are helpful. $(3)$ An incorrect patch resolves the crash but in the process modifies (or amputates) functionality. $49$ CrashFixer patches are incorrect, raising an interesting open research question of how to enable agents to verify that intended functionality is preserved after a crash is mitigated, in the absence of a practice of fine-grained unit tests in the kernel. We show examples of these categories in Section 7.
Total Crashes
Resolved
Plausible
Patches
Helpful
Patches
Incorrect
Patches
68 10 9 49
Table 3. Patch Utility using 1.5-pro-002 (with execution).
6.4. RQ3. Cost and Latency
For a single kBenchSyz bug, with $4$ trees (having $4$ nodes each), we complete a maximum of $16$ debug cycles. Assuming the localized file content is $100K$ tokens, and that CrashFixer generates $8K$ tokens in each step of the pipeline, we estimate the average cost to be \$$\sim\$13$ for each bug. This estimate assumes that the LLM is always successful at getting a compilable patch without having to repeat the hypothesis-to-patch decomposition. In practice, this isn’t true and will vary for each LLM. Each debug cycle can take a maximum of $\sim 15$ minutes. However, as each tree can run in parallel, the average bug should ideally take $\sim 1$ hour to complete in the worst case, given enough machines.
7. Case studies on CrashFixer patches
This section shows example patches generated by CrashFixer, categorized into three types: (1) plausible patches, (2) helpful but inaccurate patches, and (3) patches that resolve the crash at the cost of modifying (or amputating) functionality. They are for both open and kBenchSyz bugs.
[…]
[breaklines=true,linenos,tabsize=4]ccode_snippets/ubsan_bug.c UBSAN Shift-Out-of-Bounds Bug.
Figure 5. Hypothesis for UBSAN bug.
7.1. Plausible Patches
To study plausible patches, we scraped the Syzkaller dashboard and filtered for bugs that: (1) had their first crash reported after March 10th, 2025, (2) included a valid reproducer, and (3) remained open as of April 10th, 2025. This yielded 42 bugs, of which kGymSuite successfully reproduced approximately 30.
Because these bugs are recent and lack developer patches, we conservatively consider every file mentioned in the crash report as a localization candidate—excluding files that appear frequently across all reports. This results in an average of $\sim$13 candidate files per bug. Using this setup, we construct a dataset analogous to kBenchSyz.
While running CrashFixer on this dataset, we encountered compatibility issues between recent Syzkaller versions and the latest Linux kernel (unr, 2025), which led to unrelated kernel crashes even after the original bugs were resolved. These issues disrupted our automated pipeline. As we work to address them, we present two plausible patches generated by CrashFixer and analyze them in detail.
(1) UBSAN: shift-out-of-bounds bug. Shown in Listing 7, this bug (ubs, 2025a) was reported by UBSAN (ubs, 2025b) in the Wacom USB device devier. The Syzkaller reproducer calls wacom_parse_and_register with features->pktlen (i.e. packet length) set to 00. Function __kfifo_alloc later calculates the size of the required memory by calling roundup_pow_of_two with the value of pktlen, which unfortunately caused a “shift-out-of-bounds” error.
CrashFixer Fix. CrashFixer resolved this crash at level three of a debug tree. Its patch (lines $20$ to $29$) modifies function “wacom_devm_kfifo_alloc” to return early with an error code when pktlen is 00 (line $26$). This error is then handled by “wacom_parse_and_register” and the control goes to an error handling label on line $13$ (fail:) that returns early from this function. Figure 5 shows the CrashFixer generated hypothesis.
Developer Discussions. Concurrently to our experiment, developers discussed this bug on the kernel mailing list. From the discussion, we confirm that the CrashFixer hypothesis and patch match developer understanding. Listing 7 also shows the developers’ suggested fix (lines 44447777).
Differences. CrashFixer modifies a different function from the developer, though the error-handling logic is identical. CrashFixer returns a different error code (EINVAL) from the developer (ENODEV). Since boundaries between error codes are not strict, EINVAL is also applicable to the invalid argument.
(2) General Protection Fault bug. Shown in Listing 7.1, the function “expand_vma_in_place” tries to expand a virtual memory area (or VMA) in-place and saves this expanded vma into both vma and vrm->vma on line $3$. However, if the function “vma_merge_extend” fails, this sets both variables to NULL. When function “vrm_uncharge” is called, the dereference vrm->vma (line $12$) crashes the kernel.
[…]
[breaklines=true,linenos,tabsize=4]ccode_snippets/orig_vma_code.c General Protection Fault bug (gen, 2025).
[…]
[breaklines=true,linenos,tabsize=4]ccode_snippets/vma_fixed_code.c Fixes for the General Protection Fault bug.
Figure 6. Hypothesis for General Protection Fault bug.
CrashFixer Fix. Shown in Listing 7.1, CrashFixer breaks down the problematic line into two separate assignments (lines $3$ and $10$, respectively). Even if vma is NULL, the call to vrm_uncharge at line 6 no longer dereferences NULL. It is one of the edits at level one in a debug tree. We show CrashFixer’s associated hypothesis in Figure 6.
Developer Discussions. Developer discussions consider a likely patch in Listing 7.1 (lines 1515151521212121), which also breaks down the problematic line into two separate assignments.
Differences. CrashFixer additionally sets vrm->vma to NULL (on line $7$) before returning an error. Since vrm represents only the memory remap request, not the actual underlying memory map, and vrm->vma is not used after the error return, resetting vrm->vma has no ill effect and is defensive programming.
For examples of CrashFixer generated plausible patches in kBenchSyz refer to Appendix A.4.
[…]
[breaklines=true,linenos,tabsize=4]ccode_snippets/helpful_patch.c Helpful Patch for Warning Bug(war, 2025b).
7.2. Helpful Patches
The second class of patches are those that are helpful but not perfect. These patches allow developers to understand what “could” work to help prevent the kernel crash but might add unnecessary or remove useful code in the process. Listing 7.1 shows a kBenchSyz bug (war, 2025b) and its developer (lines 55556666) and CrashFixer (lines 1414141429292929) patches. The developer patch adds a condition to check if inode->i_size < 00 to ensure that a given symbolic link is valid. In contrast, the CrashFixer patch does something similar but is not identical. It adds two new if conditions that check for both < 00 and >= SIZE_MAX-1 on lines $18$ and $14$, respectively. However, it forgets to set the right function pointer (like line $7$) for this corner case, adds a debug statement (line $26$), and returns different error codes (lines $17$ $21$).
[…]
[breaklines=true,linenos,tabsize=4]ccode_snippets/amputation_patch.c Incorrect Patch for Warning Bug(war, 2025a).
7.3. Incorrect Patches
Finally the third class of patches are those that resolve the crash but at the expense of amputating functionality. We show an example of such a patch to a kBenchSyz bug that reaches a WARN call (war, 2025a). In Listing 7.2, CrashFixer resolves this crash by deleting the WARN_ON_ONCE call on line $2$ and merging error handling into the default case of a switch statement. In contrast, the developer patch (fix, 2025) adds a case to the switch statement, and modifies the code for future requirements.
8. Discussion
While our results are promising, designing LLM agents for complex systems like the Linux kernel remains challenging.
Semantic Correctness vs. Crash Resolution. Crash reproduction serves as a validation signal, but absence of a crash does not guarantee semantic correctness. Kernel patches must respect subtle invariants (e.g., locking, memory, interrupt safety) not captured in crash reports or as tests. Violating these can lead to latent bugs. Incorporating static checkers or symbolic reasoning post-patch is an open direction.
Incomplete Observability. Crash reports from tools like Syzkaller often lack full context—missing inputs, hardware states, and thread interleavings—forcing the agent to reason under partial observability. This hampers root-cause analysis, especially for concurrency bugs. Enhancing agent context with lightweight modeling or hardware-aware tracing may improve accuracy.
Reproducible Behavior. Agent behavior can be unstable due to limited feedback. We try to minimize this via structured reasoning and forest-based edits, but finer-grained feedback is needed for reliable control.
Filtering Bad Patches. Automatically discarding (observably) incorrect patches is a promising area. Neural, static, or runtime techniques to detect faulty edits are critical for safe system-level patching, as is incorporation of powerful patch-validation techniques such as targeted fuzzing (Tan et al., 2023; Gong et al., 2025).
Trust and Adoption. Kernel maintainers demand high-quality, explainable patches. Building trust requires LLMs to support reasoned, explainable, and interactive interactions to gain developer acceptance.
9. Conclusions
This paper introduced CrashFixer, the first LLM-based software-repair agent targeting the Linux kernel. The paper presents an evaluation of various repair strategies for kernel bugs and showcases the value of performing an explicit root-cause analysis before attempting to manipulate the complex logic of the Linux kernel. We also evaluated CrashFixer’s capabilities on still-open bugs, and we found at least two patch suggestions that were considered plausible to resolve the reported bug. Rather than closing the book on automated Linux bug repair, our work on CrashFixer and kGymSuite demonstrates that the overworked and overstressed kernel developers may have a new tool, and renewed hope, towards improving kernel reliability and security.
Appendix A Appendix
A.1. Details of Kernel Builds on kGymSuite
Figure 7. Kernel Build Time (from scratch vs from SuiteCache) on a sample kBenchSyz data point
kGym build cycles. The kGym platform performs the following steps when a user issues a kernel-build job.
(i) kGym checks if the user-specified Linux kernel tree (e.g. torvalds, net or linux-next) is available in its local Linux-repository cache. If unavailable, kGym will git pull the corresponding remote tree into this cache and clone a repository copy into its working directory.
(ii) kGym will then git checkout the given user-specified commit and attempt to apply the user patch (specified as a git diff patch). If the patch application succeeds, kGym applies additional (and necessary) “backport” commits to resolve issues that arise from incompatibilities between old kernel source code and new compilation toolchains.
(iii) Finally, kGym then invokes GNU make to build the kernel. Once the compilation is done, kGym uploads the built binaries to a cloud storage bucket.
As mentioned, the existing kernel-build process in kGym is slow as it compiles the entire kernel from scratch regardless of the number of files the patch modifies. As shown in Figure 7, the above-detailed process can take a significant amount of time to complete. As depicted, when building from scratch using kGym, a kBenchSyz bug can take $\sim 19$ minutes to build its corresponding kernel on a $8$-vCPU Google Compute Engine (c3-standard-8 GCE). This build time can be possibly reduced by running jobs on compute-heavy machines - reducing to $\sim 4$ to $5$ minutes on a powerful $88$-vCPU GCE (c3-standard-88).
Model
Strategy
Edit
Reasoning
[…]
SCS Execution
111 bugs
LCS Execution
144 bugs
1.5-pro-001 Hyp-SR then Patch-SC $\mathcal{F}$ 4 4 1 54.67 56.33
1.5-pro-001 Execution + Hyp-SR/Patch-SC $\mathcal{F}$ 4 4 1 55 (51/57) 64 (60/69)
1.5-pro-002 Hyp-SR then Patch-SC $\mathcal{F}$ 4 4 1 58.33 64.33
1.5-pro-002 Execution + Hyp-SR/Patch-SC $\mathcal{F}$ 4 4 1 64 (62/68) 71 (68/75)
Table 4. Average (min/max) crashes resolved in $3$ runs with/without execution logs. Note that the SCS/LCS subsets of kBenchSyz are those for which we could produce an execution trace, so they are not the same SCS/LCS subsets as in Table 2 of the paper.
We note that kGym adopts the above design as it serves to provide a generic building tool that can support numerous kernel versions. As each bug in kBenchSyz addresses a different kernel version (git commit), the build configurations for each bug are also unique. Hence the current design serves to provide a uniform build process applicable to all the kBenchSyz bugs ($279$ in total). However, to accurately mimic a crash-fixing workflow, the underlying platform must support ways to amortize the cost of compilation.
SuiteCache. Thus, rather than starting compilation from scratch, kGymSuite begins compilation from SuiteCache- an “intermediate cache” unique to each sample in kBenchSyz. For each sample, SuiteCache includes the source files and the pre-compiled object files for the entire Linux source code at the specified commit-id. Given that most human patches modify very few files, using SuiteCache avoids rebuilding the kernel’s unchanged files; allowing for significant speedups (Linux has more than $50k$ files and $20M$ lines of code). An example of this speedup is shown in the last column of Figure 7. As seen, even when using an $88$-vCPU GCE (c3-standard-88), the time taken to build the kernel drops from $\sim 240$ to $135$ seconds. Additionally, when reducing the core count from $88$ to $8$, the “from scratch” build time increases $\sim 4.6$x, but the SuiteCache build time increases by only $\sim 1.2$x. Thus, for the purposes of “crash resolution”, the performance of kGymSuite is practically deployment agnostic, achieving roughly the same performance despite downgrading from a $88$-vCPU to a $8$-vCPU GCE.
Figure 8. Distribution of kernel build time for kBenchSyz
To better understand the impact of using SuiteCache, we plot the histogram distribution of the $88$-vCPU build times across the entire kBenchSyz dataset in Figure 8. As shown, the distribution of kernel build times in the entire kBenchSyz dataset shifts significantly to the left when using SuiteCache. On average, we measure a meaningful speedup of 42.69% in kernel-build times using SuiteCache.
Compilation checks. kGym’s current design makes it impossible for users to issue “compilation check” jobs that specifically check if a patch is compilable. This tool proves useful if CrashFixer needs to choose a single patch from a collection of plausible patches before attempting an entire kernel-build and reproducer run. Currently, the only way to perform this with kGym is to issue multiple kernel-build jobs and to wait until the jobs succeed or fail.
To enable quick compilation checks, we revisit the kernel compilation process. Importantly, we realize that in practice, to confirm if a patch is compilable, it seems sufficient to check if we can successfully compile the files that are modified by the patch. We validate this heuristic in our use of kGymSuite by confirming that we do not observe linker failures after this pre-check. Thus, compiling other unmodified kernel files and linking objects into a kernel binary are, in practice, unnecessary steps that kGymSuite can skip. Both these optimizations can be conveniently solved with the help of SuiteCache. More specifically, when performing a compilation-check job, kGymSuite pulls the relevant SuiteCache into the working directory and can directly check if a patch is compilable by compiling solely the modified files and entirely skipping the linking process. This results in extremely quick turnaround times. In practice, kGymSuite spends most of the time downloading and extracting SuiteCache and $\sim 2$ to $3$ seconds per patch to check if the patch is compilable. As a result, kGymSuite can conveniently check $10$ patches within a span of $30$ seconds on a $88$-vCPU machine.
Thus SuiteCache is a simple yet efficient solution to both slow kernel builds and slow compilation checks.
A.2. Evaluation Result with Execution Traces
In Table 4, we report our results when running CrashFixer with execution information. For a fair comparison, we report numbers on the subset of data for which our execution collection process works ($255$/$279$ data points). This “execution” subset removes $6$ SCS bugs and $18$ LCS bugs. As shown, when using execution traces, we see relative improvements in both LCS and SCS. In LCS, we see relative improvements of $10.36\%$ for 1.5-pro-002 and $13.61\%$ for 1.5-pro-001. In SCS, we see a $9.7\%$ relative increase for 1.5-pro-002 and consistent results for 1.5-pro-001.
A.3. Memory leak bug execution traces
Patch for Null Pointer De-reference Bug (pla, 2025b)
[…]
Patch for Memory Leak Bug (pla, 2025a)
[…]
Patch for Out of Bounds Bug (pla, 2025c)
[…]
Patch for Warning Bug (pla, 2025d)
[…]
Table 5. CrashFixer generated plausible patches for different bugs in kBenchSyz.
Figure 9. Minimized execution trace for the sample memory leak bug.
In Figure 9, we depict the minimized execution traces for our example memory leak bug. As shown, “dvb_create_media_graph” is the function (in red highlights) present in the crash stack trace. The output from “stack trace anchoring” matches this function. Then the forward and backward expansions collect $4$ and $2$ functions, respectively. In particular, the execution log’s last function – ‘dvb_media_device_free()” is the cause of the memory leak. Exposing such a minimized execution trace can help CrashFixer concentrate on important and relevant executed functions.
A.4. Plausible patches for kBenchSyz bugs
In Table 5, we show examples of plausible patches for sample kBenchSyz bugs. As seen, the plausible patches span various bug types like null pointer dereferences, memory leak bugs, out-of-bound bugs and warning bugs.
A.5. Hypothesis Generation Prompt
<issue>
{CRASH TEXT}
</issue>
<code>
[start of file_1]
{file_1 text}
[end of file_1]
[start of file_2]
{file_2 text}
[end of file_2]
</code>
<execution>
Call Stack (from earliest to most recent):
{execution trace here}
</execution>
**Task**: Based on the given issue, code
snippet and **execution** please provide
a high-level reasoning for the root cause
of the issue followed by a natural language
explanation of a possible solution.
**Instructions**:
1. Write a natural language explanation of
a solution in the format ‘<solution> ...
</solution>‘.
2. Replace ‘...‘ with your solution explanation
between the ‘<solution>‘ and ‘</solution>‘ tags.
3. The solution explanation must **NOT** contain
**large** code snippets. It must be primarily a
natural language explanation with optional
references to variables and short pieces of code.
**Example**:
<code>
[start of dir_X/dir_Y/script.c]
int calc_sum(int max_num) {
    int sum = 0;
    for (int var=0; var<=max_num; var++) {
[…]
        sum -= var;
    return sum;
int main() {
    int MAX = 10;
    int sum = calc_sum(MAX);
    sum -= 3;
    printf("Sum of numbers from 1 to 10 is %d
", sum-1);
[end of dir_X/dir_Y/script.c]
</code>
<execution>
main()  [ACTIVE]
Ψ\_   calc_sum()  [executed and popped]
</execution>
**Answer of Example**:
<solution>
The snippet should calculate and print the sum
of numbers from 1 to 10. As shown in the execution
logs, the ‘main()‘ function calls a helper
function ‘calc_sum()‘ to perform the summation.
However, the variable ‘sum‘ is decremented in
the loop instead of incremented. To correct this
bug we should modify the code in the ‘for loop‘
so that the variable ‘sum‘ is incremented correctly.
Additionally, in the ‘main()‘ function, there is
an erroneous subtraction (‘sum -= 3‘) and a faulty
printf statement. To correct this bug, we should
(a) remove the extraneous subtraction and (b)
correct the printf statement to print ‘sum‘
instead of ‘sum-1‘.
</solution>
IMPORTANT NOTE: The generated hypothesis **must** be
related to the functions in the <execution> ...
</execution> tags as those functions **DEFINITELY**
contain the buggy code. So please first inspect the
functions in the <execution> ... </execution> tags
and then suggest a hypothesis that involves modifying
the code in those executed functions. Any hypothesis
that does **not** involve modifying the code in the
functions **mentioned** in the <execution> ...
</execution> tags will very likely be a **wrong**
hypothesis.
A.6. Patch Generation Prompt
<issue>
{CRASH TEXT}
</issue>
<code>
[start of file_1]
{file_1 text}
[end of file_1]
[start of file_2]
{file_2 text}
[end of file_2]
</code>
<execution>
Call Stack (from earliest to most recent):
{execution trace here}
</execution>
<solution>
{high-level reasoning prompted previously}
</solution>
**Task:** Generate a list of code modifications
based on the provided code snippets, the issue
and the natural language explanation of a
possible solution.
### Instructions:
1. **Import necessary libraries** as needed.
2. **Analyze the execution** to identify the
modifications required.
3. **Analyze the code** to identify the
modifications required.
4. **Copy and repeat** the natural language
explanation of a possible solution provided
in ‘<solution>...</solution>‘
5. **Format** the modification list as shown below:
    - Use ‘<reason>...</reason>‘ to describe
the reason for the modification.
    - Use ‘<file>...</file>‘ to specify the file path.
    - Use ‘<original>...</original>‘ to include
the original code snippet.
    - Use ‘<patched>...</patched>‘ to include
the modified code.
6. **Multiple modifications** may be provided
if necessary.
### Format for Modifications:
<solution>
</solution>
// Modification 1
<reason>
</reason>
<file>
</file>
<original>
</original>
<patched>
</patched>
// Modification 2
<reason>
</reason>
<file>
</file>
<original>
</original>
<patched>
</patched>
// Modification 3
### Example:
**Code Snippet:**
<code>
[start of dir_X/dir_Y/script.c]
int calc_sum(int max_num) {
    int sum = 0;
    for (int var=0; var<=max_num; var++) {
[…]
        sum -= var;
    return sum;
int main() {
    int MAX = 10;
    int sum = calc_sum(MAX);
    sum -= 3;
    printf("Sum of numbers from 1 to 10 is %d
", sum-1);
[end of dir_X/dir_Y/script.c]
</code>
<execution>
main()  [ACTIVE]
Ψ\_   calc_sum()  [executed and popped]
</execution>
<solution>
The snippet should calculate and print the
sum of numbers from 1 to 10. As shown in
the execution logs, the ‘main()‘ function
calls a helper function ‘calc_sum()‘ to
perform the summation. However, the variable
‘sum‘ is decremented in the loop instead of
incremented. To correct this bug we should
modify the code in the ‘for loop‘ so that
the variable ‘sum‘ is incremented correctly.
Additionally, in the ‘main()‘ function, there
is an erroneous subtraction (‘sum -= 3‘) and
a faulty printf statement. To correct this
bug, we should (a) remove the extraneous
subtraction and (b) correct the printf
statement to print ‘sum‘ instead of ‘sum-1‘.
</solution>
**Sample Answer:**
<solution>
The snippet should calculate and print the
sum of numbers from 1 to 10. As shown in
the execution logs, the ‘main()‘ function
calls a helper function ‘calc_sum()‘ to
perform the summation. However, the variable
‘sum‘ is decremented in the loop instead of
incremented. To correct this bug we should
modify the code in the ‘for loop‘ so that the
variable ‘sum‘ is incremented correctly.
Additionally, in the ‘main()‘ function, there
is an erroneous subtraction (‘sum -= 3‘) and
a faulty printf statement. To correct this
bug, we should (a) remove the extraneous
subtraction and (b) correct the printf
statement to print ‘sum‘ instead of ‘sum-1‘.
</solution>
// Modification 1
<reason>
The below change makes sure that the variable
"sum" is being incremented correctly inside
the while loop.
</reason>
<file>
dir_X/dir_Y/script.c
</file>
<original>
    for (int var=0; var<=max_num; var++) {
[…]
        sum -= var;
</original>
<patched>
    for (int var=0; var<=max_num; var++) {
[…]
        sum += var; /* changed the
        subtraction to addition */
</patched>
// Modification 2
<reason>
The below change makes sure that the extra
erroneous subtractions are removed and the
printf statement is corrected.
</reason>
<file>
dir_X/dir_Y/script.c
</file>
<original>
    sum -= 3;
    printf("Sum of numbers from 1 to 10 is %d
", sum-1);
</original>
<patched>
    printf("Sum of numbers from 1 to 10 is %d
", sum);
</patched>